OpenStudy (otonogold):

Let f(x)=x^2+6 and g(x)= x+8/x. find ( g o f)(-7) A: -55/7 B: 384/7 C: 295/49 D: 63/55

OpenStudy (anonymous):

\[(g º f)(x) \implies g(f(x))\]

OpenStudy (otonogold):

I cant see what you type out, it looks like \[(g º f)(x) \implies g(f(x))\].. my open study is acting up lol

OpenStudy (otonogold):

@agent0smith

OpenStudy (agent0smith):

Find f(-7) first. Then we'll find g(f(-7))

OpenStudy (agent0smith):

f(x)=x^2+6 f(-7) means plug in -7 for x

OpenStudy (otonogold):

-43

OpenStudy (agent0smith):

You're forgetting something, when you squared -7. f(-7)=(-7)^2+6

OpenStudy (otonogold):

55?

OpenStudy (agent0smith):

Good, now g(f(-7)) = g(55) = ?

OpenStudy (agent0smith):

g(x)= x+8/x so g(55) means plug in 55.

OpenStudy (otonogold):

3033/55 ?

OpenStudy (otonogold):

8/x a fraction or division ?

OpenStudy (agent0smith):

g(x)= x + 8/x OR g(x)= (x+8)/x which is it?

OpenStudy (otonogold):

its a fraction on my paper g(x)= (x+8)/x

OpenStudy (agent0smith):

g(55)= (55+8)/55

OpenStudy (agent0smith):

Now just simplify the numerator.